OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 2x – 5y = –14 x – 2y = –5

OpenStudy (anonymous):

x = 3, y = 4

OpenStudy (anonymous):

let x=2y-5 substitute into 2x-5y=-14 then, 2(2y-5)-5y=-14 4y-10-5y=-14 -10-y=-14 y=4 substitute into x-2y=-5 then x-2(4)=-5 x=3